### v1

# Effects of dynamical friction on perturbations for evolving dark energy

Abstract
We explore the impact of dynamical friction on cosmological scales and show its influence on the evolution of perturbations. In particular, considering smooth and clustering dark energy models, we describe the role played by friction by selecting two main hierarchical models, i.e., the first where the friction term is proportional to the Hubble rate, whereas the second where friction is induced by the dark energy pressure. The second approach generalises the first and translates the idea that pressure is a general relativistic effect, motivating why friction might arise once barotropic dark energy fluids are considered. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, using the formalism of the spherical collapse model. Whilst dynamical friction has very small effects and thus it cannot be excluded a priori, dissipative pressure friction leads to a substantial slow down in the evolution of perturbations. This can be particularly inferred within the halo mass function, for which we also employ corrections due to dark energy clustering. To this end, in order to discern detectable deviations from the standard cosmological model, we thus highlight where dissipation effects might play a significant role at large scales.
pacs:
98.80.Jk, 98.80.-k, 98.80.Es
I Introduction
Ever since the discovery of cosmic acceleration Riess et al. (1998); Perlmutter et al. (1998), several probes undoubtedly showed that the universe expansion rate is currently speeding up, through something unknown dubbed dark energy, or in the simplest case by vacuum energy Martin (2012). In order to disclose the dark energy nature, several theoretical models have been proposed throughout the last two decades Copeland et al. (2006).
Besides the existence of low- and high-redshift tests Aviles et al. (2012), alternatives can make use of morphological features of the observable universe, among which high-precision galaxy clustering and weak gravitational lensing measurements, to check how dark matter can cluster. The clustering process is intimately related to small perturbations induced by cosmic fluids in the universe and, in fact, any dark energy model might overcome successfully the aforementioned low- and high-redshift tests, together with morphological probes.
In this respect, the concordance background paradigm, i.e., the best candidate to explain how the acceleration occurs, dubbed the $\Lambda$CDM model Bull et al. (2016), where $\Lambda$ represents the cosmological constant and CDM the cold dark matter component, unequivocally appears to be the best description of the Cosmic Microwave Background (CMB) and Large Scale Structure (LSS) surveys. Phrasing it differently, the concordance approach depends on the fewest number of free parameters possible Perivolaropoulos and Skara (2022), suggesting that any dark energy model is statistically disfavoured with respect to it [^1] For the sake of clarity, the $\Lambda$CDM model requires six parameters, albeit at late-times, one single parameter, namely the mass density, approximates well the overall dynamics. Thus, to falsify the model, a compelling dark energy scenario might be described by one parameter only. The only class of dark energy models prompting one parameter only, being different from $\Lambda$CDM is the dark fluid, see e.g. Arbey (2006); Aviles and Cervantes-Cota (2011); Luongo and Muccino (2018); D’Agostino et al. (2022). Model-independent tests seem to certify the goodness of the standard background model at both low and intermediate levels Luongo and Muccino (2020). Shafieloo et al. (2018).
So, apparently, there is no reason to go further as the $\Lambda$CDM model describes the universe so well at large scales. However, even though the cosmological constant $\Lambda$, within the context of general relativity, describes perfectly well the cosmic speed up, a new dark sector physics seems to be required in order to heal current tensions Di Valentino et al. (2021); Douspis et al. (2018) and the associated caveats related to the $\Lambda$CDM model itself.
Consequently, the complete understanding of nonlinear structure formation becomes central to model the power spectrum and then to infer deviations from the concordance paradigm. In this respect, rich clusters can be used to constrain cosmological models as they represent the largest virialized objects in the universe [^2] Hence, they can be put in relation with the linear matter power spectrum, whereas their temperature can be used to infer the cluster mass..
Several alternatives to the standard concordance model have been investigated Copeland et al. (2006); Capozziello et al. (2019a), with models spanning from extending general relativity, up to barotropic dynamical dark energy, passing through interacting dark energy models [^3] Most of these approaches assume some form of energy transfer that affects both background and perturbations. These models usually fail to fit CMB data or are severely constrained..
Among all, the study of friction at cosmological scales and dynamical friction, namely the friction supposed to evolve according to the equation of state of dark energy, represent plausible features of cosmological perturbations not yet fully explored. Plausible effects induced by dissipation may compete against clustering during the overall process that induces the overdensity formation, suggesting that including friction is of crucial interest for understanding, in dark energy scenarios, how clustering is modified accordingly.
Motivated by the above points, we here investigate the effects due to dynamical friction at large scales. In this respect, we consider both smooth and clustering dark energy models, limiting the analysis, for the sake of simplicity, to models with a linearly evolving equation of state Chevallier and Polarski (2001); Linder (2003). We choose this class of models to guarantee observations to hold, as dark energy scenarios seem to favour a weakly-evolving dark energy term, instead of more complicated functions [^4] For these reasons, we exclude extended theories of gravity and we focus on general relativity only. Cosmic tests seem to agree with this assumption Calzà et al. (2020); Muccino et al. (2021). Capozziello et al. (2014); Gruber and Luongo (2014); Capozziello et al. (2015); Dunsby and Luongo (2016); Aviles et al. (2014); Capozziello et al. (2020); Aviles et al. (2013a, b); Luongo (2011); de la Cruz-Dombriz et al. (2016); Capozziello et al. (2019b); Capozziello et al. (2018); Luongo (2013); Aviles et al. (2017). In particular, to describe the role played by friction, we single out two main models. In the first scenario, friction is proportional to the Hubble rate, whereas the second shows a more complicated correction that is mostly proportional to the dark energy pressure. The corresponding effects of friction are investigated at the level of linear and nonlinear perturbations, adopting the basic demands of spherical collapse theory. The evolution of perturbations is studied and the initial conditions properly inferred. We display our numerical bounds, showing that the main discrepancy occurs for the first hierarchy, at both the levels of linear and nonlinear perturbations. Consequences on the large scale structures induced by the presence of friction are reported in detail, showing the main differences to standard frictionless cases.
Our numerical findings indicate that there is no reason to exclude a priori the presence of friction in cosmological realms. Even though the standard cosmological model does not take into account dissipation effects, we predict suitable bounds within the presence of dissipation can in principle be introduced. Since this property would significantly affect the overall clustering, we propose the viable ranges to detect it. Our numerical results suggest that there is no precise justification to exclude the presence of friction in cosmological realms a priori. Even though the standard cosmological model neglects dissipation effects, we propose potential limits of friction within clustering. To do so, given the considerable impact of this recipe in structure formation, we propose feasible ranges for its detection, with the aim to detect dissipation effects in cosmology.
The paper is organised as follows. In Sect. II we present a detailed derivation of the equations of motion for the collapsing fluid and in Sect. III we discuss the models used in this work. In Sect. IV we discuss the evolution of linear perturbations while the study of nonlinear perturbations is presented in Sect. V. In Sect. VI we show the effects of dynamical friction on the halo mass function. We conclude our work in Sect. VII.
II Extended spherical collapse model with friction
In this section we derive the expressions needed to study the evolution of linear and nonlinear perturbations of dark matter in a dark energy model where the dark energy component is smooth and only affects the background evolution. We will then extend the formalism to include perturbations also for dark energy.
The spherical collapse model assumes that perturbations are spherically symmetric and isolated Tomita (1969); Gunn and Gott (1972); Bernardeau (1994); Lee and Ng (2010). A further assumption, which allows to greatly simplify the equations, is that perturbations are described by a top-hat profile, i.e., the profile does not change in space but it is constant over the size of the perturbation. The value of the density can, therefore, only change in amplitude, but the shape of the profile stays the same. This is quite a crude approximation, but it has been proven to be very effective in describing the nonlinear evolution of perturbations, showing a good agreement between the theoretical and the numerical mass function in $N$-body simulations for both dark energy Pace et al. (2010) and modified gravity models Pace et al. (2014a).
This basic picture has been improved in several ways, mainly relaxing the assumption of spherical symmetry and isolated perturbations. Ohta et al. (2003, 2004) assumed triaxial (ellipsoidal) perturbations and took into account tidal fields; Del Popolo et al. (2013a, b, c); Pace et al. (2014b) considered the effects of shear and rotation (we will define them in a precise way later) assuming an expression derived by modelling the angular momentum of galaxies and clusters while Reischke et al. (2016); Pace et al. (2017a); Reischke et al. (2018) proposed an approach based on the Zel’dovich approximation. Finally, Pace et al. (2019); Pace and Schimd (2022) used deviations from sphericity to study the virialization of cosmic structures and Abramo et al. (2007, 2008); Chang et al. (2018) considered perturbations also in the dark energy component.
Here, we build upon all these works by further extending the formalism to include the effects of dynamical friction in $\Lambda$CDM and dark energy models. Hence, we will derive the equations of motion for matter perturbations taking into account shear, rotation and dynamical friction.
Dynamical friction manifests itself during the evolution of perturbations and affects their evolution slowing down their collapse. This means that cosmic structures require more time to form. As in Newtonian dynamics, friction affects the equations of motion and, being proportional to velocity, it modifies the Euler equation, which now reads (Carrilho et al., 2022)
$\frac{\partial\mathbf{u}}{\partial t}+(2H+\eta)\mathbf{u}+\left(\mathbf{u}\cdot\nabla_{\mathbf{x}}\right)\mathbf{u}+\frac{1}{a^{2}}\nabla_{\mathbf{x}}\phi=0\,.$ (1)
In the previous equation, $H$ is the Hubble function, $\mathbf{u}$ represents the comoving peculiar velocity defined as $\mathbf{v}=aH\mathbf{x}+a\mathbf{u}$, with $a$ the scale factor, $\mathbf{x}$ the comoving position and $\mathbf{v}$ the fluid velocity. The perturbed potential is denoted by $\phi$. Finally, $\eta=\eta(t)$ is an arbitrary function of time denoting the evolution of dynamical friction. As it is evident from Eq. (1), dynamical friction has dimensions of an inverse of time.
Continuity and Poisson equations are not affected by dynamical friction and their functional form is as usual given, respectively, by
$\,\delta^{\prime}+(1+\delta)\tilde{\theta}=0\,,$ (2a)
$\,\nabla^{2}\phi=\frac{3}{2}a^{2}H^{2}\Omega_{\rm m}(a)\delta\,,$ (2b)
with $\Omega_{\rm m}(a)$ the matter density parameter, $\delta=\delta\rho/\bar{\rho}$ the fluid density perturbation and the prime represents the derivative with respect to $\ln{a}$.
We can combine the continuity, Euler and Poisson equations into a single equation describing the evolution of the perturbation $\delta$. Assuming as independent variable $\ln{a}$, the full nonlinear equation is
$\delta^{\prime\prime}+$ $\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\delta^{\prime}-\frac{4}{3}\frac{{\delta^{\prime}}^{2}}{1+\delta}-\frac{3}{2}\Omega_{\rm m}(a)\delta(1+\delta)-$
$\qquad\left(\tilde{\sigma}^{2}-\tilde{\omega}^{2}\right)(1+\delta)=0\,,$ (3)
where $H^{2}\tilde{\sigma}^{2}=\sigma^{2}=\sigma_{ij}\sigma^{ij}$ and $H^{2}\tilde{\omega}^{2}=\omega^{2}=\omega_{ij}\omega^{ij}$ are the shear and rotation tensors, respectively. They are defined as
$\sigma_{ij}=\frac{1}{2}(\partial_{i}u^{j}+\partial_{j}u^{i})-\frac{1}{3}\theta\delta_{ij}\,,\quad\omega_{ij}=\frac{1}{2}(\partial_{i}u^{j}-\partial_{j}u^{i})\,,$ (4)
with $\theta=\nabla\cdot\mathbf{u}$ the divergence of the peculiar velocity and $\delta_{ij}$ the Kronecker’s delta symbol.
If we linearise Eq. (II), we obtain the modified equation describing the evolution of the growth factor:
$\delta^{\prime\prime}+\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\delta^{\prime}-\frac{3}{2}\Omega_{\rm m}(a)\delta=0\,,$ (5)
Both the linear and nonlinear expressions assume the same form of those obtained for dark scattering models Baldi and Simpson (2017); Carrilho et al. (2021, 2022). In these models there is an interaction between the cold dark matter and the dark energy component which gives rise to an exchange of energy between the two fluids. The equations of motion lead to an additional friction term which modifies the coefficient of $\delta^{\prime}$ and depends on the cross section and mass of the dark matter particles. Here we want to move a step further and study the physical consequences of different implementations of the dynamical friction on the nonlinear evolution of perturbations, in particular on the mass function.
The formalism is easily extended to include the effects of clustering dark energy. To focus on this, let us now consider the following perturbation equations for the dark energy component (Abramo et al., 2007, 2008; Abramo et al., 2009a, b; Pace et al., 2017b):
$\delta_{\mathrm{de}}^{\prime}+3(s_{\rm eff}-w_{\mathrm{de}})\delta_{\mathrm{de}}+$
[…]
$\tilde{\theta}_{\mathrm{de}}^{\prime}+\left(2+\frac{H^{\prime}}{H}\right)\tilde{\theta}_{\mathrm{de}}+\frac{\tilde{\theta}_{\mathrm{de}}^{2}}{3}+$
[…]
where the prime implies again the derivative with respect to $\ln{a}$. Clearly, $\tilde{\theta}_{\mathrm{de}}=\theta_{\mathrm{de}}/H$ and $w_{\mathrm{de}}$ is the dark energy equation of state at the background with the corresponding sound speed defined by $c_{\rm eff}^{2}=c^{2}s_{\rm eff}=\delta P_{\mathrm{de}}/\delta\rho_{\mathrm{de}}$, where the speed of light, $c$, appears for dimensional reasons. This quantity is zero for the standard $\Lambda$CDM model and for classes of dark fluids, while it is fixed to $s_{\rm eff}=1$ for Quintessence. The different values that this can acquire have huge consequences in cosmology: a small sound speed implies structures to form at several scales, while high sound speeds correspond to stiff matter, suggesting perturbations to be strongly suppressed, i.e., involving dark energy that affects the background only. Consequently, it is licit to presume $s_{\rm eff}\ll 1$, guaranteeing that dark energy perturbations become relatively important in analogy to matter perturbations, albeit always subdominant (Batista and Pace, 2013). Note that, in general, $\theta_{\rm de}\neq\theta_{\rm m}$, as friction only affects (cold dark) matter.
II.1 A comment on dynamical friction
At this point, a deeper discussion of the results obtained is necessary. It is well known that an alternative way to discuss cosmological perturbation is based on the evolution of the physical radius of the perturbation $R$. In fact, many authors already considered the effect of dynamical friction from this point of view. The evolution of the (effective) radius of the perturbation $R$, including angular momentum $L$ and dynamical friction $\eta$, in a $\Lambda$CDM model is given by Lahav et al. (1991); Bartlett and Silk (1993); Peebles (1993); Antonuccio-Delogu and Colafrancesco (1994); Del Popolo and Gambera (1998, 1999) [^5] In these references, only single parts of the full expression are investigated or reported.
$\frac{\mathrm{d}^{2}R}{\mathrm{d}t^{2}}=-\eta\frac{\mathrm{d}R}{\mathrm{d}t}-\frac{GM_{\rm m}}{R^{2}}+\frac{\Lambda c^{2}}{3}R+\frac{L^{2}}{M_{\rm m}R^{2}}\,,$ (7)
where $t$ is the cosmic time, $c$ the speed of light, $G$ Newton’s constant, $\Lambda$ the value of the cosmological constant and $M_{\rm m}$ the mass of the dark matter perturbation. This expression has been explicitly written down in Del Popolo et al. (2019) to derive a mass-temperature relation for clusters in $\Lambda$CDM and modified gravity models.
Following Pace et al. (2019), we can easily extend the previous expression to an arbitrary dark energy model described by an equation-of-state parameter $w_{\rm de}$ and express the angular momentum contribution in terms of the shear $\sigma$ and rotation $\omega$ term
$\frac{\mathrm{d}^{2}R}{\mathrm{d}t^{2}}=-\eta\frac{\mathrm{d}R}{\mathrm{d}t}-\frac{GM_{\rm m}}{R^{2}}-\frac{GM_{\rm bck,de}}{R^{2}}(1+3w_{\rm de})-\frac{\sigma^{2}-\omega^{2}}{3}R\,,$ (8)
where we defined
$M_{\rm m}=\frac{4\pi}{3}\bar{\rho}_{\rm m}(1+\delta)R^{3}\,,\quad M_{\rm bck,de}=\frac{4\pi}{3}\bar{\rho}_{\rm de}R^{3}\,.$ (9)
In the previous expressions, $\bar{\rho}_{\rm m}$ and $\bar{\rho}_{\rm de}$ are the background matter and dark energy density, respectively and $M_{\rm bck,de}$ represents the amount of dark energy included in a sphere of radius $R$. This expression can be further generalised to take into account dark energy perturbations as in Pace and Schimd (2022)
$\frac{\mathrm{d}^{2}R}{\mathrm{d}t^{2}}=$ $\,-\eta\frac{\mathrm{d}R}{\mathrm{d}t}-\frac{GM_{\mathrm{m}}}{R^{2}}-\frac{GM_{\mathrm{de}}}{R^{2}}(1+3s_{\rm eff})$
$\,-\frac{GM_{\rm bck,de}}{R^{2}}(1+3w_{\mathrm{de}})-\frac{\sigma^{2}-\omega^{2}}{3}R\,,$ (10)
where
$M_{\mathrm{de}}=\frac{4\pi}{3}\bar{\rho}_{\mathrm{de}}(a)\delta_{\mathrm{de}}R^{3}\,,$ (11)
is the mass associated to dark energy perturbation. Note that this equation does not say anything whether also dark energy experiences dynamical friction.
As shown in Appendix B, Eq. (8) leads to
$\delta^{\prime\prime}+$ $\,\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\delta^{\prime}-\frac{4}{3}\frac{{\delta^{\prime}}^{2}}{1+\delta}-\frac{3}{2}\Omega_{\rm m}(t)\delta(1+\delta)-$
$\left(\tilde{\sigma}^{2}-\tilde{\omega}^{2}\right)(1+\delta)-\frac{3\eta}{H}(1+\delta)=0\,,$ (12)
which contains the two additional terms $-\tfrac{3\eta}{H}(1+\delta)$ and $\frac{\eta}{H}\delta^{\prime}$. The latter is a proper friction term. At the linear level, it reads $-\tfrac{3\eta}{H}$ and leads to conceptual problems. It acts as a source term such that the particular solution is, for a model where $\eta=\eta_{0}H$ and for an EdS model, $\delta=-2\eta_{0}$. This implies that for $a\to 0$ an overdensity originates from an underdensity, or said in other terms, primordial perturbations are small, but do not go to zero as expected.
As noted also in Carrilho et al. (2022), this is due to the fact that dynamical friction only depends on the peculiar velocity of dark matter and that while for the overdensity we use comoving coordinates, for the evolution of the radius we use physical coordinates. In particular the radius of the perturbation is the physical radius. When we go from the evolution of the density perturbation $\delta$ to that of $R$, using the same procedure outlined in Appendix B, the equation for the latter is (see also Carrilho et al., 2022)
$\frac{\mathrm{d}^{2}R}{\mathrm{d}t^{2}}=$ $\,-\frac{GM_{\rm m}}{R^{2}}-\frac{GM_{\rm bck,de}}{R^{2}}(1+3w_{\rm de})$
$\,-\frac{\sigma^{2}-\omega^{2}}{3}R-\eta\left(\frac{\mathrm{d}R}{\mathrm{d}t}-HR\right)\,.$ (13)
This equation contains two terms originating from friction: one contribution dampens the collapse ($-\eta\,\mathrm{d}R/\mathrm{d}t$) and the other ($\eta HR$) enhances it. It is this term, $\eta HR$, which exactly cancels the term $3\eta(1+\delta)/H$ in Eq. (II.1) and thus curing the aforementioned problem, i.e. that perturbations are sourced, rather than being slowed down by friction.
This can be also seen by defining the comoving variable $y\equiv\tfrac{R}{R_{i}}\tfrac{a_{i}}{a}=\lambda\tfrac{R}{a}$, so that Eq. (II.1) becomes
$\ddot{y}+(2H+\eta)\dot{y}+\frac{1}{2}H^{2}\Omega_{\rm m}(a)\delta y+\frac{\sigma^{2}-\omega^{2}}{3}y=0\,,$ (14)
with $\delta$ the density contrast of the perturbation defined as
$\delta=[1+\delta_{i}]\left(\frac{a_{i}}{a}y\right)^{-3}-1\,.$ (15)
If we instead consider the original formulation of dynamical friction via the equation of motion of the physical radius $R$, Eq. (8), we arrive to
$\ddot{y}+(2H+\eta)\dot{y}+\frac{1}{2}H^{2}\Omega_{\rm m}(a)\delta y+\eta Hy+\frac{\sigma^{2}-\omega^{2}}{3}y=0\,,$ (16)
where, once again, we see a term arising from the mixture of comoving and physical quantities. With respect to Carrilho et al. (2022) we also took into account shear and rotation terms. From now on, we will neglect their contribution in our analysis and set them to zero, effectively considering an exact spherical symmetry, as we want to focus on the effects of friction and avoid an interplay among them and friction. The effects of shear and rotation on structure formation have been studied before for both smooth and clustering dark energy models and for different implementations, as already discussed in Sect. II. We will not delve into them again, but we refer the reader to Del Popolo et al. (2013a, b, c); Pace et al. (2014b); Reischke et al. (2016); Pace et al. (2017a); Reischke et al. (2018); Pace et al. (2019); Pace and Schimd (2022) for details.
In Sect. III we detail how to calculate the time dependence of $\eta$ with respect to time. Here we simply summarise our findings. For the dynamical friction model of Antonuccio-Delogu and Colafrancesco (1994) we obtain $\eta=\eta_{0}\,H$ with $\eta_{0}$ constant and for dissipative pressure friction we find $\eta=\eta_{0}(\eta_{1}+\eta_{2}|P/P_{0}|^{\Gamma})\,H_{0}$, with $\eta_{1}$, $\eta_{2}$ and $\Gamma$ all constant.
III Theoretical models of friction
In this Section we describe the two models investigated in this work.
III.1 Evolution of dynamical friction
Here, we follow closely the works by Antonuccio-Delogu and Colafrancesco (1994) and Del Popolo (2009), to which we refer for more details.
In a scenario with hierarchical structure formation, structures of size $R$ form around the local maxima of the primordial density field smoothed over a scale of size $R$ Bardeen et al. (1986); Bond (1988); Colafrancesco et al. (1989). Assuming an initial Gaussian density field, the comoving averaged number density of peaks with central height $\nu\equiv\delta(z)/\sigma(M,z)$, with $\delta$ the overdensity and $\sigma$ the variance of the field, is given by
$\mathcal{N}_{\rm pk}(\nu|R)\,\mathbf{d}\nu=\frac{1}{(2\pi)^{2}R_{\ast}^{3}}G(\nu,\gamma\nu)\,e^{-\nu^{2}/2}\,,$ (17)
where the spectral parameters are $R_{\ast}=\sqrt{3}\sigma_{1}/\sigma_{2}$ and $\gamma=\sigma_{1}^{2}/(\sigma_{0}\sigma_{2})$, the function $G$ is given in Eq. (4.4) of Bardeen et al. (1986) and $\sigma_{l}$, with $l\in\{0,1,2\}$, are the spectral moments defined as
$\sigma_{l}^{2}(R,z)=\frac{1}{2\pi^{2}}\int\,\mathrm{d}k\,k^{2(1+l)}P(k,z)W(kR)\,,$ (18)
and $W(kR)$ is a filtering function.
In this system, a test particle experiences a gravitational field which consists of a part associated to the smooth global mass distribution and a stochastic one which originates from the particle number fluctuations and giving rise to a frictional force $-\eta\mathbf{v}$ where $\eta$ is the dynamical friction coefficient and $\mathbf{v}$ the macroscopic velocity.
Following Antonuccio-Delogu and Colafrancesco (1994), one can show that the peaks of the local density field with central height $\nu\geq\nu_{\rm c}$, with $\nu_{\rm c}$ a critical threshold, contribute to the dynamical friction. The number of these objects can be calculated under the condition that the peak radius $r_{\rm pk}(\nu\geq\nu_{\rm c})$ is negligible with respect to the average peak separation $n_{\rm a}(\nu\geq\nu_{\rm c})^{-1/3}$.
The values for $\nu_{\rm c}$ and the spectral moments will depend on the particular cosmological model considered, and the precise value of the spectral moments depend on the filtering scale $R$. Here we do not pretend to derive exact values for the quantities above for every individual model we investigate in this work. Our aim is merely to estimate the order of magnitude of the relevant quantities, in particular the dynamical friction coefficient $\eta$ and understand how this parameter impacts the evolution of linear and nonlinear structures in $\Lambda$CDM and dark energy models.
For a standard cosmological model with flat CDM spectrum Bond and Efstathiou (1984), Antonuccio-Delogu and Colafrancesco (1994); Del Popolo (2009) found $\nu_{\rm c}\approx 1.24$ and for a peak of mass $\approx 10^{9}\,h^{-1}\,M_{\odot}$ one obtains $\gamma\approx 0.442$ and $R_{\ast}\approx 45.2\,h^{-1}\,\text{kpc}$.
Using the virial theorem, it is rather straightforward to show that the dynamical friction coefficient reads
$\eta=$ $\,\frac{4.44(Gm_{\rm a}n_{\rm a})^{1/2}}{N_{\rm tot}}\log{[1.12N_{\rm tot}^{2/3}]}$
$=$  $\,4.44\sqrt{\frac{3\Delta}{8\pi}}\frac{\log{[1.12N_{\rm tot}^{2/3}]}}{N_{\rm tot}}H=\eta_{0}\,H\,,$ (19)
where we used the fact that $m_{\rm a}n_{\rm a}=\rho_{\rm m}=\bar{\rho}_{\rm m}\Delta$, with $\Delta$ the average overdensity of the perturbation. In principle, $\Delta$ would depend on the virialization recipe and on the cosmological model. However, here for simplicity, we will assume it to be constant, of the order of $\mathcal{O}(100)$.
This result shows that $\eta\propto H$ and that the value of $\eta_{0}$ depends on the total number of peaks $N$ inside a protocluster, decreasing when $N$ increases. As discussed in Antonuccio-Delogu and Colafrancesco (1994), $N$ is considered to be approximately constant for clusters. For $N_{\rm tot}=10^{3}$ we find $\eta_{0}\approx 3\times 10^{-2}$, while for $N_{\rm tot}=10^{4}$ we have $\eta_{0}\approx 4\times 10^{-3}$. Regardless of the exact value of $\eta_{0}$, the main conclusion is that $\eta_{0}\ll 1$.
Note also that this result is not in contrast with previous works, such as Antonuccio-Delogu and Colafrancesco (1994), as there an EdS model was assumed. Indeed, for such a model, neglecting radiation, we have $H\propto a^{-3/2}$. In fact, Antonuccio-Delogu and Atrio-Barandela (1992) showed that it is possible to associate a relaxation timescale with dynamical friction in galaxy clusters and this is of the order of the Hubble time, i.e., $1/\eta\propto H^{-1}$.
Finally, rather than expressing $\eta$ in terms of the total number of peaks $N$, it is worth to use the mass of the perturbation $M$. Recalling that $M=m_{\rm a}N$, it is straightforward to obtain
$\eta=4.44\sqrt{\frac{3\Delta}{8\pi}}\frac{\log{[1.12(M/m_{\rm a})^{2/3}]}}{(M/m_{\rm a})}H=\eta_{0}\,H\,,$ (20)
where we assume $m_{\rm a}=10^{9}\,h^{-1}\,M_{\odot}$. It is evident that $\eta_{0}$ is smaller for massive objects rather than low-mass perturbations.
Dynamical friction is a consequence of the fact that general relativity considers a pressure term that is absent in Newtonian dynamics. Dynamics is influenced by pressure and it is natural to follow this direction, i.e., friction originated by pressure to extend the standard dynamical friction model. We will describe this approach in the next section.
III.2 Dissipative pressure friction
Standard fluids can be grouped into two main classes: Newtonian and non-Newtonian. Naively, the latter is found when the friction coefficient, $\eta$, depends upon the variation of shear velocity with respect to the radial coordinate. Our approach is simplified by virtue of the cosmological principle, i.e., the velocity field does not depend upon radial coordinate. In this scenario, however, our friction coefficient is a function of the Hubble rate only. In other words, we do not emphasise a time-dependence of it, assuming the simplest stationary case. It is clearly possible to generalise this case, including the effects due to the dynamics of the universe. Following the standard treatment, one can assume that the friction coefficient is a function of temperature and/or of thermodynamic quantities such as pressure, entropy, enthalpy and so forth. In the simplest case, we can assume what follows
$\eta=\left(\eta_{1}+\eta_{2}\left|\frac{P}{P_{0}}\right|^{\Gamma}\right)H_{0}\,,$ (21)
where $\eta_{1},\eta_{2}$ and $\Gamma$ are free real constants, with $P$ the total pressure of the universe and $P_{0}$ the total pressure today. The overall $H_{0}$ factor is just a normalization factor which makes the two coefficients $\eta_{1}$ and $\eta_{2}$ dimensionless. From the second Friedmann equation,
$\frac{\ddot{a}}{a}=-\frac{4\pi G}{3}\left(\rho+\frac{3P}{c^{2}}\right)\,,$ (22)
where $\rho$ and $P$ represent the total density and pressure and remembering that $\frac{\ddot{a}}{a}\equiv\dot{H}+H^{2}$, combining Eqs. (21) and (22), we infer
$\eta=\left[\eta_{1}+\eta_{2}\left|\frac{2\dot{H}+3H^{2}}{2\dot{H}_{0}+3H_{0}^{2}}\right|^{\Gamma}\right]H_{0}\,.$ (23)
At this stage, we distinguish two cases, as below reported.
    Slow-rolling approximation, namely $3H^{2}\gg 2\dot{H}$, leading to
    $\eta=\left[\eta_{1}+\eta_{2}\left|\frac{H}{H_{0}}\right|^{2\Gamma}\right]H_{0}\,,$ (24)
    reducing to our previous case, discussed above, with $\Gamma=\tfrac{1}{2}$, $\eta_{1}=0$ and $\eta_{2}=\eta_{0}$. This certifies that our previous approach turns out to be a suitable approximation that works well in a phase in which $H^{2}$ dominates over the dynamics of the universe, albeit it does not account for the whole stages of its evolution.
    Fast-rolling approximation, namely $3H^{2}\ll 2\dot{H}$, having
    $\eta=\left[\eta_{1}+\eta_{2}\left|\frac{\dot{H}}{\dot{H}_{0}}\right|^{\Gamma}\right]H_{0}\,,$ (25)
    that shows a quite different case with respect to the one previously developed, where the dynamical effects, $\sim\ddot{a}$, are taken into account explicitly.
In our computations, we investigate the consequences of Eq. (21), prompting the main differences between the two limiting cases featured in Eqs. (24) and (25). To this end, it appears essential to work out different sets of $\Gamma$ values that are not known a priori.
Assuming that friction increases at large $z$, we require [^6] Remember that $\frac{\mathrm{d}}{\mathrm{d}x}|f(x)|=\frac{f(x)}{|f(x)|}\frac{\mathrm{d}f(x)}{\mathrm{d}x}$.
$\frac{\mathrm{d}\eta}{\mathrm{d}t}=\eta_{2}\,\Gamma\,H_{0}\,\left|\frac{P}{P_{0}}\right|^{\Gamma-2}\frac{P\dot{P}}{P_{0}^{2}}>0\,,$ (26)
and then to determine the sign of $\dot{P}$, we can invoke Eq. (22), having $\dot{P}\propto-2(\ddot{H}+3H\dot{H})$ that can be rewritten as
$\dot{P}=-2H^{3}(j-1)\frac{c^{2}}{8\pi G}\,,$ (27)
recasting in terms of the jerk parameter, defined as $j\equiv-H^{-1}\dot{q}+2q^{2}+q$ Dunsby and Luongo (2016), with $q\equiv-1-\frac{\dot{H}}{H^{2}}$, the deceleration parameter.
Thus, since $\dot{H}>0$ and $j\geq 0$ Luongo et al. (2016); Capozziello et al. (2022), we have
$\eta_{2}\Gamma<0\,,$ (28)
that guarantees that $\dot{\eta}>0$, albeit it manifestly implies a degeneracy on the signs of these free constants. Indeed, since $\frac{dz}{dt}=-(1+z)H(z)$, we have
$\eta^{\prime}\simeq$  $\,2\tilde{\eta}_{0}(1+z)^{3}P_{\rm rad}^{\Gamma-1}\,,$ (29a)
[…]
respectively, for an Einstein-de Sitter universe dominated by radiation and matter, where $\tilde{\eta}_{0}\equiv 2H_{0}\eta_{2}\Gamma$ and $\eta^{\prime}\equiv\frac{\mathrm{d}\eta}{\mathrm{d}z}$. Consequently, in view of the above results, we can take $\eta_{2}<0$ and $\Gamma>0$ or vice versa, having moreover
$\eta_{1}=$  $\,0\,,\,\textrm{and},\,\eta_{2}\in[0,1]\,,\Gamma\in[-2;-1]\,,\qquad{\rm or}$ (30a)
$\eta_{1}=$  $\,0\,,\,\textrm{and},\,\eta_{2}\in[-1,0]\,,\Gamma\in[1;2]\,,$ (30b)
where $\eta_{1}$ is arbitrarily imposed to vanish as it appears unessential in determining $\eta$ and the values ±$\pm 1$ selected for normalizing $\eta_{2}$ and $\Gamma$.
Having the above, the case $\Gamma=0$ corresponds to a constant friction term, while the case $\eta_{2}=0$ implies a frictionless model. Consequently, we argue that
    the first case described above implies a model where friction acts attractively, i.e., the contribution acts as a standard fluid within the Zeldovich limit. Hence the total friction, taking also into account the cosmological expansion is increased,
    the second case described above implies a model where friction acts repulsively, i.e., the sign is compatible with the dark energy equation of state. We will not consider this case here as it does not represent the ordinary behaviour of friction on matter.
From the aforementioned considerations, we fit our models within the intervals of priors determined above. We discuss our findings in what follows.
IV Evolution of linear perturbations
In this Section we discuss the evolution of linear perturbations and show how they are affected by dynamical friction.
Even though small, $\eta$ affects the initial conditions of Eq. (5), similarly to what happens in early dark energy and massive neutrinos scenarios. This happens because the friction term $\eta/H$ is small but not negligible. To elucidate this concept, let us denote with $\eta_{\rm ini}=\eta(a_{\rm ini})/H(a_{\rm ini})\ll 1$ the value to be used at initial time to evaluate the initial conditions. In general, at early time, a given cosmological model is well approximated by an EdS model, so that we can assume $\Omega_{\rm m}(t)\approx 1$ and $H^{\prime}/H\approx-3/2$ and that the growth factor follows a power-law behaviour, $\delta=Aa^{n}$, where $A$ is an arbitrary constant amplitude which can be safely set to one. Plugging the Ansatz for $\delta$ into Eq. (5), we need to solve a second order algebraic equation for the exponent $n$:
$n^{2}+\left(\frac{1}{2}+\eta_{\rm ini}\right)n-\frac{3}{2}=0\,,$ (31)
whose growing mode solution is, for $\eta_{\rm ini}\ll 1$, $n_{+}=1-2\eta_{\rm ini}/5$. This result is expected from physically considerations: friction slows down the collapse and $n_{+}<1$ when $\eta\neq 0$. We can compare this result with a generic early dark energy model. If we denote with $\Omega_{\rm e}\ll 1$ the amount of early dark energy, we have $\Omega_{\rm m}\approx 1-\Omega_{\rm e}$ and we find for the growing mode $n_{+}=1-3\Omega_{\rm e}/5$. This same expression applies to massive neutrinos, if we replace $\Omega_{\rm e}$ with the fraction of massive neutrinos $f_{\nu}=\Omega_{\nu}/\Omega_{\rm m}$, with $\Omega_{\nu}$ the neutrinos density parameter. Note the similarity of the different physical systems. Also in this case, therefore, structure formation is slowed down. This means that to match the pace of structures in the $\Lambda$CDM model, the initial amplitude of perturbations must be higher. We will discuss this in more detail when investigating the evolution of nonlinear perturbations.
After this discussion, we can now study the evolution of the growth factor. We consider two clustering species, baryons (denoted with the subscript $\mathrm{b}$) and dark matter (denoted with the subscript $\mathrm{c}$), of which only the latter is affected by dynamical friction. The two equations to be solved are
[…]
and we define the total growth factor as
$\delta_{\rm m}=\frac{\Omega_{\rm b}\delta_{\rm b}+\Omega_{\rm c}\delta_{\rm c}}{\Omega_{\rm b}+\Omega_{\rm c}}\,.$ (33)
IV.1 Initial settings
If we consider a single clustering species, we can assume that at early times $\delta=A\,a^{n}$, where both $A$ and $n$ are constant. This leads to a second-order algebraic equation for $n$ whose solution, for $\eta=\eta_{0}H$ and $\eta_{0}\ll 1$, is $n=1-\tfrac{2}{5}\eta_{0}$. We can proceed in a similar way if we consider both dark matter and baryons. Also in this case we can assume a power-law, such that $\delta_{\rm b}=A\,a^{n}$ and $\delta_{\rm c}=B\,a^{n}$, with $A$, $B$ and $n$ constants to determine. We assume that the two perturbations have the same slope but different amplitudes. If we plug this Ansatz in the previous equations, we obtain, taking into account that at early times the background is extremely well represented by an EdS model,
[…]
From the first equation we can solve for the Poisson term and plug it into the second equation, in order to find a relation between the initial perturbations
$\frac{\delta_{\rm b}}{\delta_{\rm c}}=1+\frac{\eta_{\rm ini}}{n+1/2}\,,$ (35)
which shows that the baryon perturbation amplitude is increased with respect to that of cold dark matter, as baryons are not affected by friction. Also note that for vanishing friction, $\delta_{\rm b}=\delta_{\rm c}=\delta_{\rm m}$, as expected.
Figure 1: Evolution of the growth factor for the $\Lambda$CDM model for the two recipes used in this work to describe cosmic friction. Top panel: dynamical friction as parametrised in Eq. (20). Bottom panel: dissipative pressure friction as parametrised in Eq. (21). Different colours and line styles refer to different values of the free parameters.
Plugging Eq. (35) into the equation for $\delta_{\rm c}$, leads to an equation for the exponent $n$:
$n^{2}+\left(\frac{1}{2}+\eta_{\rm ini}\right)n-\frac{3}{2}\left(1+\Omega_{\rm b}\frac{\eta_{\rm ini}}{n+1/2}\right)=0\,.$ (36)
Note that, at early times $\Omega_{\rm b}=\Omega_{\rm b,0}/\Omega_{\rm m,0}$, with $\Omega_{\rm m,0}=\Omega_{\rm b,0}+\Omega_{\rm c,0}$.
Figure 2: Evolution of the growth factor for the six dark energy models investigated in this work with dynamical friction. Line styles and colours are as in Fig. 1.
Even though it is possible to rewrite it as a third order equation, it is simpler to solve it as it is, to avoid going through complex radicals. We can, however, use the cubic equation to investigate the behaviour of the initial slope when $\eta_{\rm ini}\ll 1$. The equation to be solved is
$n^{3}+(1+\eta_{\rm ini})n^{2}-\frac{1}{4}(5-2\eta_{\rm ini})n-\frac{3}{4}\left(1+2\Omega_{\rm b}\eta_{\rm ini}\right)=0\,.$ (37)
If $\eta_{\rm ini}=0$, it is easy to see that the solutions are $n=-\tfrac{1}{2}$, $n=-\tfrac{3}{2}$ and $n=1$, of which only the latter two are the physical ones, corresponding to the decaying and growing mode, respectively. At first order in $\eta_{\rm ini}$ we obtain $n=1-\tfrac{2}{5}\eta_{\rm ini}(1-\Omega_{\rm b})$, where it is understood that $\Omega_{\rm b}$ is evaluated at early times. The term in parenthesis is positive for realistic values of the baryon density, hence, also in this case we can see that friction dampens structure formation. If we set, for simplicity $1-\Omega_{\rm b}\approx 1$, we recover the previous result, showing the excellent agreement between the two approaches.
Since the two equations of motion are different, we can not combine them into a single equation for the variable $\delta_{\rm m}$. Therefore, to highlight the importance of friction and the differences induced by having two species rather than one, we consider the evolution of the linear growth factor $D_{+}$ as a function of redshift $z$ considering either one or two species. As differences among the two different cases are of the order of $0.1\%$, for simplicity, we will consider just the case where both dark matter and baryons are affected by dynamical friction. This will also simplify the discussion when considering clustering dark energy models.
Figure 3: As in Fig. 2, but with dissipative pressure friction.
IV.2 Growth factor
In this Section we present the effects of friction on the evolution of the linear growth factor, using the two recipes described above. We first consider its effect on the $\Lambda$CDM reference model. We then compare it to six different dark energy models, two of them characterised by a constant equation of state, $w_{\rm de}=-0.9$ and $w_{\rm de}=-1.1$ (DE1 and DE2), and four with a linearly time-dependent equation of state, $w_{\rm de}=w_{0}+w_{a}(1-a)$, with $w_{0}$ and $w_{a}$ constant (Chevallier and Polarski, 2001; Linder, 2003). We choose the following combinations of parameters $(w_{0},w_{a})$: $(-0.9,\pm 0.3)$ (DE3 and DE4) and $(-1.1,\pm 0.3)$ (DE5 and DE6).
In Fig. 1 we present the evolution of the linear growth factor, divided by the scale factor, for the $\Lambda$CDM model, showing results for both dynamical friction (upper panel) and dissipative pressure friction (bottom panel). The considerations which we are going to make for this model will apply also to the other models studied in this work.
As $\eta$ opposes the formation of structures, perturbations grow more slowly with respect to a model where the friction term is absent. Hence, because we want the same rate today (all the curves have been normalised to unity today), structures need to growth faster in the past. This explains the differences observed. Note that this behaviour is analogous to $\Lambda$CDM versus EdS. Therefore, the higher is the value of $\eta$, the higher the difference from a model without friction.
For dynamical friction, in general, differences are very small and only appreciable for the highest value chosen, $\eta_{0}=10^{-2}$. Effects of dissipative pressure friction follow the same trend. Dissipative pressure friction behaves similarly to dynamical friction and when the models are all normalised to one today, they present an increase in the growth factor. These considerations explain the behaviour found in Fig. 1. We also note that changing $\Gamma$ has no effect on the evolution of perturbations, but this is only true for the $\Lambda$CDM model, where density, and hence pressure, do not evolve in time. Finally, the effects induced by dissipative pressure friction are much larger than for dynamical friction. This happens because whilst the contribution of dynamical friction is constant in time, dissipative friction changes with time and its effects are stronger later in the cosmic evolution.
Similar considerations can be made for Figs. 2 and 3 where we show the effects of dynamical friction for the first and dissipative friction for the latter for the six dark energy models investigated in this work. We note that the effect of dynamical friction for these models is qualitatively the same as for the $\Lambda$CDM model when we consider the relative difference with respect to a model without friction. This is clearly understood as dark energy is only affecting the background and the dynamical equation to be solved is the same for both $\Lambda$CDM and the dynamical dark energy models. Furthermore, the correction due to $\eta_{0}$ is constant in time.
A deeper discussion about the effect of dissipative pressure friction is worth. In this case, because pressure changes in time, the parameter $\Gamma$ has an effect and the higher it is (in absolute value), the higher is the difference with respect to the model not affected by friction. As for the $\Lambda$CDM model, a positive $\eta_{2}$ implies a higher growth factor. In addition, exactly because pressure changes in time, the observed effect depends on the particular equation of state. Models 1, 3, 4 and 5 share similar effects for the cases $(+1,-2)$ and $(+1,-1)$ while for model 2 differences start to increase. Model 6 presents the strongest effect for the combination $(+1,-2)$ and this is due to the fact that the equation of state stays in the phantom regime all the time and, as a consequence, pressure is higher (in absolute value).
V Evolution of nonlinear perturbations
The study of nonlinear perturbations is based on the approach of the spherical collapse, whose equations have been derived in Section II. We consider once again two perturbed fluids, dark matter and baryons, where only the first experiences the effects of dynamical friction. To derive the initial conditions of the full nonlinear equation we consider that the averaged matter density $\delta_{\rm m}$ diverges Pace et al. (2010). However, for numerical stability, we do not use $\delta_{\rm m}$, but rather its inverse, so that the collapse takes place when $1/\delta_{\rm m}\to 0$, as discussed in Pace et al. (2017b).
Some important quantities which can be defined within the formalism of the spherical collapse are the turn-around scale factor, $a_{\rm ta}$ and the nonlinear overdensity at turn-around $\zeta$. The first represents the time at which the spherical perturbation reaches its maximum extension before collapsing and shrinking, while the latter is defined as $\zeta=1+\delta_{\rm NL}(a_{\rm ta})$. The nonlinear overdensity at turn-around is constant for an Einstein-de Sitter model but for any other model, including the $\Lambda$CDM one, changes in time. We verified that dynamical friction has either a negligible or very small effect both on the turn-around scale factor $a_{\rm ta}$ or on the nonlinear overdensity at turn-around, $\zeta$. When considering dissipative pressure friction, instead, the value of $a_{\rm ta}$ changes up to $2\%$, while $\zeta$ differs up to $15\%$ with respect to a model without friction.
Figure 4: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
Two additional and very important quantities, also from an observational point of view, are the linearly extrapolated overdensity $\delta_{\rm c}$ and the virial overdensity $\Delta_{\rm vir}$. The linear overdensity $\delta_{\rm c}$ is, together with the linear growth factor, the main ingredient of the halo mass function which, as the name says, represents the linear evolution of matter density perturbations, whose initial conditions are found to have the nonlinear matter density evolution to diverge at collapse time. The linear overdensity thus represents the solution of a linear differential equation but with initial conditions derived from its nonlinear counterpart. Its time evolution as a function of the collapse redshift $z$ for the $\Lambda$CDM model is presented in Fig. 4. We show the results for the dark energy models in Section A in Figs. 10 (dynamical friction) and 11 (dissipative pressure friction), to which we refer for the discussion of our findings.
Even though the overdensity $\delta_{\rm c}$ is a linear quantity, the differential equation is evolved starting from nonlinear initial conditions and this explains why differences are enhanced with respect to the growth factor but are reduced with respect to the nonlinear overdensity $\zeta$. Focusing again on the $\Lambda$CDM model, for the dynamical friction case, we see that, as expected, differences are enhanced when the free parameter $\eta_{0}$ increases, so that for $\eta_{0}=0.01$ the linear overdensity is about 0.2% lower than when $\eta_{0}=0$, regardless of the collapse redshift considered. The highest difference is at $z=0$ and at the higher redshift all the models recover the corresponding value for the EdS model. When friction becomes stronger, $\delta_{\rm c}$ is lower, even though the initial conditions are enhanced. This is because to collapse at the same time of a model without friction, perturbations require a lower threshold for the collapse to compensate the presence of friction.
In the bottom panel of Fig. 4 we consider dissipative pressure friction applied to the $\Lambda$CDM model. As expected, the effects induced by friction are higher than for dynamical friction and, in analogy to the growth factor, the parameter $\Gamma$ is irrelevant (but it will not be for dynamical dark energy models). Differences, now, are of the order of $-5\%$ and $\delta_{\rm c}$ decreases in agreement to what found above. This time, though, for all the combinations of parameters, we recover the standard $\Lambda$CDM model at higher redshift. This is expected as for dissipative pressure friction the pressure terms grows with time and is, therefore, important at relatively late times, when the dark energy component dominates.
The virial overdensity $\Delta_{\rm vir}$ represents the value the nonlinear overdensity reaches at virialization. To determine how virialization proceeds, we follow Carrilho et al. (2022) for interacting dark energy models. Even though this is not the physical case we are studying here, for smooth dark energy, the equations of motion are the same and dark matter experiences, in both cases, an additional dynamical friction force.
If we denote with $K$ the total kinetic energy of the system and with $W$ its potential energy, the expression for the virial theorem is modified as
$2\langle K\rangle+\langle W\rangle+\sum_{i}\langle\mathbf{F}_{i}^{\rm fric}\cdot\mathbf{r}_{i}\rangle=0\,,$ (38)
where $\mathbf{F}_{i}^{\rm fric}$ is a non-conservative force, $\mathbf{r}_{i}$ the physical position of particles over which the summation is extended to. The friction force can be written as
$\mathbf{F}_{i}^{\rm fric}=-m\eta\,x_{i}^{\prime}\,,$ (39)
with $m$ the mass of individual particles.
We can make explicit the different terms entering into the virial theorem. To do so, we follow the notation of Cataneo et al. (2019); Carrilho et al. (2022) and denote with [^7] We corrected a typo in the relation between $\delta$ and $y$ in Eq. (28) of Carrilho et al. (2022) and Eq. (34) of Cataneo et al. (2019).
$E_{0}=\frac{3}{10}M_{\rm m}(H_{0}\,R_{i})^{2}\,,\qquad\delta=(1+\delta_{i})\,y^{-3}-1\,,$ (40)
so that
$K=$ $\,\frac{H^{2}}{H_{0}^{2}}\left[\frac{a}{a_{i}}(\partial_{N}y+y)\right]^{2}\,E_{0}\,,$ (41a)
$W_{\rm N}=$ $\,-\Omega_{\rm m,0}\frac{1}{a\,a_{i}^{2}}\,y^{2}(1+\delta)\,E_{0}\,,$ (41b)
$W_{\rm de}=$ $\,-\frac{H^{2}}{H_{0}^{2}}(1+3w_{\rm de})\Omega_{\rm de}(a)\left(\frac{a}{a_{i}}\right)^{2}y^{2}\,E_{0}\,,$ (41c)
$W_{\rm fr}=$ $\,-2\frac{\eta}{H}\frac{H^{2}}{H_{0}^{2}}\left(\frac{a}{a_{i}}\right)^{2}y\,\partial_{N}y\,E_{0}\,,$ (41d)
with $N\equiv\ln{a}$ where we decomposed the total potential energy $W$ in its contributions, i.e., $W=W_{\rm N}+W_{\rm de}+W_{\rm fr}$, with $W_{\rm N}$ the Newtonian potential due to matter perturbations, $W_{\rm de}$ the background dark energy contribution and $W_{\rm fr}$ the friction contribution. In a more familiar form, expressing these quantities in term of the physical radius $R$, we have
$K=$ $\,\frac{3}{10}M_{\rm m}\dot{R}^{2}\,,$ (42a)
$W_{\rm N}=$ $\,-\frac{3}{5}\frac{GM_{\rm m}^{2}}{R}\,,$ (42b)
$W_{\rm de}=$ $\,-\frac{3}{5}\frac{GM_{\rm m}M_{\rm bck,de}}{R}(1+3w_{\rm de})\,,$ (42c)
$W_{\rm fr}=$ $\,-\frac{3}{5}\eta\,M_{\rm m}R\left(\dot{R}-HR\right)\,.$ (42d)
Note how in the definition of $W_{\rm fr}$, where we use the physical radius, there appears once again the term $\dot{R}-HR$ which was discussed above.
Figure 5: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for the dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
Therefore, according to Eq. (38), to find the virilization epoch, we solve the equation
$2K+W_{\rm N}+W_{\rm de}+W_{\rm fr}=0\,,$ (43)
and the virial overdensity is defined as
$\Delta_{\rm vir}=(1+\delta_{\rm NL})\left(\frac{a_{\rm c}}{a_{\rm vir}}\right)^{3}\Omega_{\rm m}(a_{\rm vir})\,,$ (44)
with $a_{\rm c}$ and $a_{\rm vir}$ the collapse and virialization scale factor, respectively. The nonlinear overdensity $\delta_{\rm NL}$ is evaluated at the virialization epoch.
Knowing the virialization epoch, using the formalism outlined above, we can also determine the ratio of the virial radius to the turn-around radius, $R_{\rm vir}/R_{\rm ta}$. This quantity gives an idea of how much the perturbation shrinks from its maximum size once it reaches the virial equilibrium. This quantity is, therefore, affected by friction but it is also cosmology- and time-dependent. It is constant in time only for an EdS model, for which $R_{\rm vir}/R_{\rm ta}=1/2$ (in absence of friction contributions). Once again, for dynamical friction differences with respect to the standard case are extremely small, but can reach a few percent in the case of dissipative pressure friction.
We can now discuss the effects of friction on the evolution of the virial overdensity $\Delta_{\rm vir}$, as defined in Eq. (44) and presented in Fig. 5 for the $\Lambda$CDM model. In the top (bottom) panel we consider dynamical (dissipative pressure) friction. Once more, the first case leads to very small effects, except for $\eta_{0}=0.01$, with differences of about $0.5\%$, relatively constant in time, with a small increase at late times when the cosmological constant dominates. For the case of dissipative pressure friction, effects are, as extensively discussed before, relevant, reaching up to $50\%$ today. These results are in perfect agreement with our previous analysis of the different nonlinear quantities, in particular of the overdensity at the turn-around $\zeta$.
Concluding, dynamical friction appears having only small signatures on the models we investigated. On the other side, albeit theoretically-motivated, the dynamical pressure model shows highly-performing changes into the shapes that are expected once friction is involved. Thus, on the one hand, the latter clearly appears more probable to be measured. Conversely, the first case appears hard to be disentangled from the background dark energy. Remarkably, both are not excluded, since it would be possible to experimentally constrain the associated free parameters to check which values are thus expected to correctly reproduce a friction contribution. For the sake of completeness, we expect the dynamical pressure model to behave less deeply, being constructed in the simplest way in which the pressure itself is modified as due to the inclusion of a friction term.
VI Effects of the modified evolution of the mass function
In this section we build upon results of previous sections and we investigate the halo mass function. This is an important observable, used to probe cosmology at late times by counting observed structures in the Universe. Differently from other observables, such as the CMB or the matter power spectrum, it is not model-independent, in the sense that an observed halo mass function relies on an assumed cosmology via the evaluation of the cosmic volume. This cosmology-dependence may be alleviated or eliminated considering observations at $z\simeq 0$.
From a theoretical point of view, the functional form of the halo mass function can be inferred from fits to $N$-body simulations (Jenkins et al., 2001; Tinker et al., 2008; Watson et al., 2013), even though analytical derivations exist (Press and Schechter, 1974; Sheth et al., 2001). In this work we make a very conservative choice and use the Sheth-Tormen (ST) halo mass function (Sheth et al., 2001). This choice is dictated by the fact that Creminelli et al. (2010) used this formulation to describe the halo mass function in clustering dark energy models with null sound speed. We will come back to this point later.
The functional form of the halo mass function for Sheth et al. (2001) is
$f(\nu)=A\sqrt{\frac{2a}{\pi}}\left[1+(a\nu)^{-p}\right]\nu\exp{\left(-\frac{1}{2}a\nu^{2}\right)}\,,$ (45)
where $\nu=\delta_{\rm c}/\sigma$, with $\sigma$ the variance of the linear matter power spectrum and $A=0.3222$, $a=0.75$ and $p=0.3$.
The mass variance of the smoothed density field is given by
$\sigma^{2}(M,z)=\frac{1}{2\pi^{2}}\,\int_{0}^{\infty}\,k^{2}\,P(k,z)\,W(kR)\,\mathrm{d}k\,,$ (46)
where the window function $W(kR)=3[\sin{(kR)}-(kR)\cos{(kR)}]/(kR)^{3}$ is the Fourier transform of the top-hat profile and the radius is associated to the halo mass via the relation $M=4\pi\bar{\rho}_{\rm m,0}R^{3}/3$ and $\bar{\rho}_{\rm m,0}$ today’s average matter density.
We obtain the linear matter power spectrum from a suitably modified version of the Einstein-Boltzmann code class [^8] https://github.com/lesgourg/class_public/ Blas et al. (2011) [^9] For simplicity, we perform calculations in the Newtonian gauge.. For internal consistency, the code always evolves the perturbations of every species, therefore, to distinguish between smooth and clustering dark energy we set cs2_fld, the variable representing $c_{\rm s}^{2}$, to $1$ and 00, respectively.
Figure 6: Evolution of the matter power spectrum $P(k)$ at $z=0$ for the $\Lambda$CDM model. In the top (bottom) panel we show results for dynamical (dissipative pressure) friction. Line styles and colours are as in Fig. 1.
Before showing our results on the halo mass function, we present in Fig. 6 the matter power spectrum for the $\Lambda$CDM model at $z=0$ only.
Unsurprisingly, dynamical friction is negligible on all the scales considered, except for the case with the largest $\eta_{0}$ and on very small scales. In this case, differences range between $5\%$ and $10\%$. Dissipative friction, instead, as already discussed before, shows a stronger effects, almost constant along the scales shown and with differences of about 25% with respect to the case without friction. As expected, differences point towards a lack of power, sign of a slowing down in the evolution of dark matter perturbations. This is due to the fact that we considered that all models have the same initial conditions, in particular the primordial amplitude of perturbations $A_{\rm s}$. This translates into a lower value of the normalization of the matter power spectrum today, $\sigma_{8}$.
For dark energy models we find qualitative and quantitative similar results, therefore we will not show additional plots. Dynamical friction has a very small effect, while dissipative pressure friction is substantially more relevant. The exact quantitative effect depends on the particular dark energy model, but its behaviour is totally analogous to what seen for the growth factor, presented in Figs. 2 and 3. As perturbations in dark energy models are very small, even when $c_{\rm s}^{2}=0$, we do not see appreciable differences between smooth and clustering dark energy models.
We can now discuss the effects of friction on the halo mass function. We present our results for the number of halos above a given mass for the $\Lambda$CDM model in Fig. 7 and for smooth dark energy models in Figs. 8 and 9.
Figure 7: Number of objects above a given mass $M$ for the $\Lambda$CDM model at $z=0$. The top panel shows the effects of dynamical friction while the bottom panel refers to dissipative pressure friction. Line styles and colours are as in Fig. 1.
Dynamical friction has a small effect on the overall mass function and becomes appreciable only for the most extreme case with $\eta_{0}=0.01$. In this case, for masses of the order of $7\times 10^{14}\,M_{\odot}$, we notice a decrease of about 20% in the cumulative number of objects. On the same scales, the effect of dissipative pressure friction is about three times higher. As we evaluated the halo mass function today, this is a consequence of the lack of power highlighted in the matter power spectrum. We also remind the reader that for the $\Lambda$CDM model, the effect of $\Gamma$ is irrelevant. This result strengthens the previous discussion on the possibility of putting tight constraints on these models.
Figure 8: Number of objects above a given mass threshold $M$ for the dark energy models studied in this work. We considered dynamical friction. Line styles and colours are as in Fig. 1.
Figure 9: As in Fig. 8, but for dissipative pressure friction.
For the smooth case, all the models have the same qualitative behaviour when considering dynamical friction and the differences with respect to a frictionless model are largely independent of the particular equation of state used. As expected, differences increase at higher masses and for all the cases here considered we observe a decrease in the number of objects which form. This is coherent with the fact that friction dampens perturbations. The higher is $\eta_{0}$, the larger are the differences. For $\eta_{0}=10^{-4}$, deviations from the standard case where dynamical friction is absent, are, once again, negligible. Differences become appreciable for $\eta_{0}=10^{-3}$ and reach a value of about 2%–3% for masses of the order of $10^{15}\,M_{\odot}\,h^{-1}$, to reach 20%–25% for $\eta_{0}=10^{-2}$.
More interesting is the case of dissipative pressure friction. Differences are enhanced with respect to the case with $\eta_{2}=0$ and can reach up to 75% of less objects. A particular case is model 6, whose equation of state is much smaller than $-1$. For this model, the suppression of structures is substantial, of the order of $50\%$ already for objects above $10^{13}\,M_{\odot}\,h^{-1}$.
The analysis shows another interesting point. Whilst the behaviour of dynamical friction is coherent among the models, and percentage differences are roughly the same, for dissipative pressure friction this is not the case any more. First of all the exact difference strongly depends on the background equation of state. In fact, excluding model 6, we observe that differences range from 50% to 75% for masses of $10^{15}\,M_{\odot}\,h^{-1}$. But more importantly, analogously to what happens for the growth factor, we see that the two cases with $\Gamma=-1$ and $\Gamma=-2$ have a different behaviour with respect to each other: in fact, for models with a quintessence behaviour (models 1 and 3), the model with $\Gamma=-2$ predicts more objects than when $\Gamma=-1$. Opposite situation when the model is always or at least for some time in the phantom regime. Thus, the lower is the equation of state, the stronger are the differences.
VII Outlooks
In this paper, we investigated the effects of friction on large scale structures. In this respect, we focused on both smooth and clustering dark energy models and limited our analysis to general relativity and to models with a linearly evolving equation of state, as these are more in line with observational data.
Specifically, to describe the role of friction, we examined two main backgrounds. The first considers friction proportional to the Hubble rate, while the second involves a more complex correction primarily related to the dark energy pressure, i.e., involving relativistic effects due to the pressure of dark energy.
We analysed the effects of friction on both linear and nonlinear perturbation regimes, employing the foundational principles of spherical collapse theory. To do so, we studied the evolution of perturbations and we determined the appropriate initial conditions to develop such perturbations.
Our findings revealed that the most significant discrepancies occur in the second hierarchical model, as it affects severely both linear and nonlinear perturbations. In this respect, we detailed the effects and consequences of friction on large-scale structures, highlighting the key differences from standard cases that do not incorporate such effects from the outset.
Further, we developed the consequences of the modified evolution of the cumulative mass function into the two aforementioned schemes. Consequences and discussions towards the underlying effects are reported, showing that our expectations are in line with current comprehension of our data. We demonstrated that the first recipe, dynamical friction proportional to the Hubble function, affects the mass function by 5% at most, while for the second model differences can be up to 75% for the range of masses probed in this work.
We also considered how perturbations of the dark energy affect the halo mass function. It turns out that the effects of clustering of dark energy is significantly more important than friction and the contribution of perturbations are strongly dependant on the background equation of state.
Consequently, with the severe exception of the second background treatment, our results seemed to suggest that there is no a priori reason to exclude the presence of friction, indicating that dissipation can potentially have a significant limiting effect on clustering.
For future works, we intend to extend our effort by involving more complicated friction models and also investigating more deeply the consequences on structure formation. Additional studies will be performed in showing how dark energy clustering can also be influenced by dark energy and/or we will clarify whether a possible interaction between the dark sector and friction may occur.
Acknowledgements
F.P. acknowledges partial support from the INFN grant InDark, the Departments of Excellence grant L.232/2016 of the Italian Ministry of Education, University and Research (MUR) and from the European Union – Next Generation EU. FP also acknowledges the FCT project with ref. number PTDC/FIS-AST/0054/2021. F.P. also acknowledges the FCT project with ref. number PTDC/FIS-AST/0054/2021. The work of O.L. is partially financed by the Ministry of Education and Science of the Republic of Kazakhstan, Grant: IRN AP19680128.
Appendix A Dark energy
In the next two subsections we will present results for the dark energy models considered in this work. We consider both smooth and clustering dark energy, but we will show figures about smooth dark energy models only, as the effect of clustering dark energy is very small and qualitatively we obtain the same results in both cases.
Figure 10: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for the dynamical friction. Line styles and colours are as in Fig. 1.
Figure 11: As in Fig. 10, but for the dissipative pressure friction.
A.1 Smooth dark energy
For smooth dark energy, we can infer analogous conclusions to the $\Lambda$CDM model. The effects of dynamical friction on the linearly extrapolated overdensity parameter $\delta_{\rm c}$ (Fig. 10) are small and of the same order of magnitude of the $\Lambda$CDM model and deviations from the case with $\eta_{0}=0$ are largely independent of the particular model considered.
For dissipative pressure friction (Fig. 11), as for the $\Lambda$CDM model, the overall amplitude of $\delta_{\rm c}$ is lower than when friction is not considered. The strength of the effect depends on the particular dark energy model. Models 2, 5 and 6 show the strongest deviations, while for the other three models differences are less pronounced. Differently from before, now the particular value of $\Gamma$ affects the final behaviour of the model. In fact, for the phantom model 6, when $\Gamma=-2$, the effect of friction is so large that the initial overdensity needs to reach unrealistic values so that perturbations can collapse. For this reason we do not show the results obtained. An indication of this comes from the bottom right panel of Fig. 3, where we see substantial differences with the standard case also at the level of the growth factor.
Regarding the effects of friction on the evolution of the virial overdensity $\Delta_{\rm vir}$, the results obtained are in line with what found for $\Lambda$CDM. In Fig. 12 we present our results for dynamical friction and in Fig. 13 for dissipative pressure friction. Also for dark energy models, dynamical friction has a very modest impact with differences of the same order of magnitude found for the $\Lambda$CDM model. Dissipative pressure friction has obviously a much stronger effect and we see an increase of the value of the virial overdensity. For models 2 and 5, in fact, the virial overdensity first increases and then decreases. It is also interesting to see that for model 3, the virial overdensity is almost constant over all the redshifts investigated.
Figure 12: Evolution of the virial overdensity $\Delta_{\rm vir}$ for the six dark energy models investigated with dynamical friction. Line styles and colours are as in Fig. 1.
Figure 13: As in Fig. 12, but for dissipative pressure friction.
Figure 14: Number of objects above a given mass threshold $M$ for the dark energy models studied in this work, when dark energy is perturbed. We considered dynamical friction. Line styles and colours are as in Fig. 1.
Figure 15: As in Fig. 14, but for dissipative pressure friction.
A.2 Clustering dark energy
In this section we discuss results for clustering dark energy models, where we allow for perturbations of dark energy. Their strength is related to the value of the effective sound speed, defined, as previously discussed, by $c^{2}_{\rm eff}=\delta P_{\rm de}/\delta\rho_{\rm de}$. To fix it, one would need to consider a scalar field and its Lagrangian. We will, therefore, consider it as a free parameter. Quintessence models have $c^{2}_{\rm eff}=1$ and their perturbations are negligible on the scales of interest Unnikrishnan et al. (2008). For $k$-essence models we can achieve smaller values and in this work we will consider $c^{2}_{\rm eff}=0$ for which dark energy perturbations are larger and the dark energy fluctuations behave like matter.
Once again, we will consider the linear (represented by the growth factor) and the nonlinear evolution (represented by $\delta_{\rm c}$ e $\Delta_{\rm vir}$), but we will not present any figure, as there are no interesting nor special features even in this case.
In the linear regime, dark energy perturbations are always very subdominant with respect to matter perturbations and the evolution of $\delta_{\rm m}$ is totally unaffected by them. Similar results are obtained when considering the evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ and of the virial overdensity $\Delta_{\rm vir}$. We did not find appreciable differences with respect to the smooth case.
When dark energy is clustering, there is always the question of what is the correct definition of the density perturbation to be used. So far we have assumed that it is enough to consider dark energy affecting matter perturbations via modification of the Poisson equation and we found that dark energy perturbations do not affect matter perturbations appreciably. One could also assume that what really matters is not matter density perturbations, but rather the total density perturbation, defined as:
$\delta(a)=\delta_{\rm m}(a)+\frac{\Omega_{\rm de}(a)}{\Omega_{\rm m}(a)}\delta_{\rm de}(a)\,,$ (47)
where it is understood that perturbations are either linear or nonlinear according to the particular quantity investigated. Here, we will limit our analysis to the evolution of linear perturbations, in particular of $\delta_{\rm c,tot}$.
Our results are presented in Fig. 16 for the case of dynamical friction and in Fig. 17 for dissipative pressure friction. We can now appreciate that the behaviour of $\delta_{\rm c,tot}$ is rather different from before and it is strongly dependant on the particular parameterisation of the equation of state used. Considering the effect of dynamical friction, we observe for models 2, 5 and 6 the typical behaviour obtained before. The linearly extrapolated overdensity approaches the EdS value (dark energy perturbations are negligible and furthermore scaled by $\Omega_{\rm de}$) and decreases at late times. For the other models, instead, we recover the EdS value at early time but at late times $\delta_{\rm c,tot}$ grows, showing that dark energy perturbations can have an appreciable effect. It is also interesting to notice that the effective value of dynamical friction becomes rather irrelevant, as all the models behave similarly to the case of no friction with $\eta_{0}=0$.
More complex is the analysis for dissipative pressure friction. Results do not only depend on the equation of state used, but also on the particular value of the parameter $\Gamma$. We thus observe a similar trend to what found in Fig. (16). Models 2, 5 and 6 behave as one would expect for smooth dark energy or for associating $\delta_{\rm c}$ to matter only. The other three models, instead, have an EdS-like behaviour at early times and grow, rather than decrease, at late times. It is also interesting to see that the dependence on the exponent $\Gamma$ becomes very shallow.
Figure 16: Evolution of the total linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for dynamical friction. Line styles and colours are as in Fig. 1.
Figure 17: Evolution of the linearly extrapolated overdensity $\delta_{\rm c}$ for the wCDM models studied here for dissipative pressure friction. Line styles and colours are as in Fig. 1.
This discussion highlights an important point about the correct definition of the density perturbations, as, even though $\delta_{\rm de}$ is smaller than $\delta_{\rm m}$ in general and have limited effects on the evolution of the latter, it can substantially affect our observables if we use $\delta_{\rm tot}$ rather than $\delta_{\rm m}$, as it was also demonstrated, discussing virialization, in Pace and Schimd (2022).
A.3 Corrections to the halo mass function
When dark energy clusters, also the halo mass function is modified. To take into accounts these modifications, we follow the work by Creminelli et al. (2010). According to the authors, one can scale the halo mass function calculated for a smooth model with the Sheth-Tormen Sheth et al. (2001) formulation with the halo mass function calculated with the Press-Schechter Press and Schechter (1974) formalism.
The halo mass function for smooth models is thus Creminelli et al. (2010)
$\frac{\mathrm{d}n}{\mathrm{d}\ln{M}}=\frac{\mathrm{d}n_{\rm{ST,s=1}}}{\mathrm{d}\ln{M}}\frac{\mathrm{d}n_{\rm{PS,s=0}}/\mathrm{d}\ln{M}}{\mathrm{d}n_{\rm{PS,s=1}}/\mathrm{d}\ln{M}}\,,$ (48)
where it is understood that when the particular dark energy is clustering, one needs to use the appropriate matter power spectrum and linear overdensity parameter $\delta_{\rm c}$. The reason for this choice is that the ratio of the Press-Schechter functions will mitigate its shortcomings. We present our results in Figs. 14 and 15 for dynamical and dissipative pressure models, respectively. We consider the ration between the clustering model with the corresponding smooth model.
For dynamical friction, we notice that the dominant contribution is that of clustering. Whether we predict more or less objects depends on the model considered. If the model has a quintessence equation of state, then clustering predicts more objects, with differences ranging from subpercent level for models 1 and 5 to a couple of percent for model 3 for masses of the order $10^{13}\,M_{\odot}\,h^{-1}$. At higher masses differences can be more than 10%. We notice that these substantial differences appear for model 3, whose equation of state is the most distant from that of the $\Lambda$CDM model. For phantom models, we observe a decrease in structure formation, general of a few percent and more substantial for model 6, as it is the model more different from $\Lambda$CDM.
Very similar results, from both a qualitative and a quantitative point of view, are obtained for dissipative pressure friction. Also in this case the clustering is the dominating effect and difference are maximised for models 3 and 6. Except for the latter, models with different values of $\Gamma$ behave very similarly to each other.
We also notice that our results are in good qualitative agreement with the original work of Creminelli et al. (2010).
Appendix B Derivation of the equation for the radius $R$
In this section we discuss how we can recover Eq. (8) from the fluid approach, but first we start from the equation of the radius $R$ and derive a second order differential equation for the matter overdensity $\delta$. We limit ourselves to the case of smooth dark energy, as that of clustering dark energy is totally analogous and will simply modify the Poisson equation with a term proportional to $\delta_{\rm de}$. To this purpose, we use the relation Engineer et al. (2000); Pace et al. (2019); Pace and Schimd (2022)
$1+\delta=\lambda\frac{a^{3}}{R^{3}}\,,$ (49)
with $\lambda$ constant and $a$ the scale factor.
Starting from Eq. (8), we need an expression for the velocity and the acceleration, which read, respectively,
$\frac{\dot{R}}{R}=$ $\,\frac{1}{3}\left[3H-\frac{\dot{\delta}}{1+\delta}\right]\,,$
$\frac{\ddot{R}}{R}=$ $\,\frac{1}{9}\left[3H-\frac{\dot{\delta}}{1+\delta}\right]^{2}+\frac{1}{3}\left[3\dot{H}+\frac{\dot{\delta}^{2}}{(1+\delta)^{2}}-\frac{\ddot{\delta}}{1+\delta}\right]\,.$
We will also use the Friedmann equation $H^{2}+\dot{H}=-\tfrac{4\pi G}{3}[\bar{\rho}_{\rm m}+\bar{\rho}_{\rm de}(1+3w_{\rm de})]$ with $G$ Newton’s gravitational constant.
With these equations, after a few algebraic steps one can easily find Eq. (II.1), which, as extensively discussed in the main body of this work, leads to problems.
We want, now, to derive the equation for $\delta$ using the fluid approach, by suitably modifying the Euler equation. This will allow us to highlight where the problems we discussed above originate.
Following Antonuccio-Delogu and Colafrancesco (1994), we use the following expression for the Euler equation [^10] Differently from Antonuccio-Delogu and Colafrancesco (1994), though, we include also nonlinear terms.
$\left.\frac{\partial\mathbf{v}}{\partial t}\right|_{\mathbf{r}}+\left(\mathbf{v}\cdot\nabla_{\mathbf{r}}\right)\mathbf{v}+\nabla_{\mathbf{r}}\Phi+\eta\mathbf{v}=0\,,$ (50)
which describes the macroscopic velocity of a fluid as in the Langevin approximation for the Brownian motion. In Eq. (50), derivatives are done with respect to the physical coordinate $\mathbf{r}$. The gravitation potential $\Phi$ consists of two terms: $\Phi_{0}$ related to a uniform matter distribution and such that $\nabla_{\mathbf{x}}\Phi_{0}=-a\ddot{a}\mathbf{x}$ and $\phi$ associated to inhomogeneities.
We now move from the time derivative with fixed physical coordinate $\mathbf{r}$ to time derivative with fixed comoving coordinate $\mathbf{x}$ using the well known relation
$\left.\frac{\partial}{\partial t}\right|_{\mathbf{r}}=\left.\frac{\partial}{\partial t}\right|_{\mathbf{x}}-H\mathbf{x}\cdot\nabla_{\mathbf{x}}\,,$
and write $\mathbf{r}=a\mathbf{x}$ and express the velocity $\mathbf{v}$ in terms of the comoving peculiar velocity $\mathbf{u}$: $\mathbf{v}=aH\mathbf{x}+a\mathbf{u}$.
Inserting these expressions, the Euler equation for the peculiar velocity is now
$\left.\frac{\partial\mathbf{u}}{\partial t}\right|_{\mathbf{x}}+(2H+\eta)\mathbf{u}+\eta H\mathbf{x}+\left(\mathbf{u}\cdot\nabla_{\mathbf{x}}\right)\mathbf{u}+\frac{1}{a^{2}}\nabla_{\mathbf{x}}\phi=0\,.$ (51)
Note how this equation contains an additional term, $\eta H\mathbf{x}$, which only depends on background quantities. It is exactly this term the source of the unphysical behaviour of the growth factor we discussed above. It is now clear to see how this term originates from using physical coordinates rather than comoving coordinates. Also, the above derivation implies that dynamical friction affects equally background and perturbations, which is though not the case.
From now on, we omit, for brevity, the subscript $\mathbf{x}$ and we imply that derivatives are done with respect to the comoving coordinate.
Taking the divergence of Eq. (51) we arrive to
$\dot{\theta}+\left(2H+\eta\right)\theta+3H\eta+\frac{1}{3}\theta^{2}+\sigma^{2}-\omega^{2}+\frac{1}{a^{2}}\nabla^{2}\phi=0\,,$ (52)
where $\theta=\nabla_{\mathbf{x}}\cdot\mathbf{u}$ is the divergence of the peculiar velocity, $\sigma$ the shear and $\omega$ the rotation.
The last step in our derivation is to use $\ln{a}$ as the independent variable and to write the Euler equation for the dimensionless variable $\tilde{\theta}=\theta/H$:
$\tilde{\theta}^{\prime}+\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\tilde{\theta}+\frac{1}{3}\tilde{\theta}^{2}+\tilde{\sigma}^{2}-\tilde{\omega}^{2}+3\frac{\eta}{H}+\frac{\nabla^{2}\phi}{a^{2}H^{2}}=0\,.$ (53)
It is easy to see that combining the continuity, Euler and Poisson equations we recover Eq. (II.1).
Dynamical friction thus enters in the equation of motion for perturbations as an additional source of friction in the Euler equation, with the result of modifying the coefficient of $\delta^{\prime}$ and adding an additional term to the overall system. This implies that also the linearised version of Eq. (II.1), which represents the growth factor, is modified and friction, therefore, also affects linear dynamics. The modified growth factor equation reads
$\delta^{\prime\prime}+\left(2+\frac{\eta}{H}+\frac{H^{\prime}}{H}\right)\delta^{\prime}-\frac{3}{2}\Omega_{\rm m}(t)\delta-\frac{3\eta}{H}=0\,.$ (54)